### v1

# humancompatible.detect: a Python Toolkit for Detecting Bias in AI Models

Abstract
There is a strong recent emphasis on trustworthy AI. In particular, international regulations, such as the AI Act, demand that AI practitioners measure data quality on the input and estimate bias on the output of high-risk AI systems. However, there are many challenges involved, including scalability (MMD) and computability (Wasserstein-1) issues of traditional methods for estimating distances on measure spaces. Here, we present humancompatible.detect, a toolkit for bias detection that addresses these challenges. It incorporates two newly developed methods to detect and evaluate bias: maximum subgroup discrepancy (MSD) and subsampled $\ell_{\infty}$ distances. It has an easy-to-use API documented with multiple examples. humancompatible.detect is licensed under the Apache License, Version 2.0.
Keywords: bias detection, toolkit
1 Introduction
With the increase in practical applications depending on machine-learning models and AI systems, new ethical problems have appeared. One particular problem is the proliferation (or even strengthening) of societal bias. Infamous examples of AI bias include hiring algorithms (Dastin, 2018) or recidivism predictors (Julia Angwin et al., 2016). In response, more and more regulatory frameworks (e.g., Regulation 2024/1689(2024), EU) and related standards (e.g., IEEE Standards Association, 2025; Schwartz et al., 2022) require testing for bias, without any clear agreement on the precise definition of said bias. Indeed, there are many bias measures (Mehrabi et al., 2021).
The lack of a single definition of bias is understandable, considering the wide range of applications of AI systems. Non-exhaustively:
  • Fairness in rankings: the main related case being job recruitment. See, for instance, Pitoura et al. (2021); Kliachkin et al. (2024); Fageot et al. (2025).
  • Fairness in recommender systems: such as the ones widely used currently by, e.g., YouTube or Spotify (Pitoura et al., 2021).
  • Spatial aspects of fairness: In contrast to continuous attributes, such as age and income, the case of discrimination based on location demands special attention (Sacharidis et al., 2023; Kyriakopoulos et al., 2025).
  • Fairness for forecasting: Consider the case of the well-known COMPAS dataset, traditionally addressed as a classification task. A forecasting-based approach, however, obtains the present-best results on the dataset (Zhou et al., 2023).
  • Fairness in sharing-economy business models: such as the model of companies such as Uber and Cabify (Zhou et al., 2024).
Even in the best-studied case of fairness of classification, most of the proposed bias measures, such as statistical parity or equality of opportunity, focus on protected groups in isolation. This approach often misses the intersections of groups or requires one to sift through an exponential number of group intersections (subgroups). Considering all subgroups is crucial, as illustrated by Figure 1, where data is marginally fair, but intersectionally biased. Another view of bias evaluation in cases when many subgroups need to be considered is bias detection. That is, detecting where bias exists in the data, rather than measuring it entirely over all protected features. Detecting such a biased subgroup allows focusing on the particular subgroup (e.g., Sharma et al., 2024). This way, one can address the most significant biases individually through mitigation techniques tailored to the particular group, instead of treating all biases the same.
We introduce humancompatible.detect [^1] Freely available at https://github.com/humancompatible/detect., an open-source toolkit of methods for bias detection. Designed for ease of use, its simple API and guided examples are accessible to non-expert users. The toolkit allows for reliable detection of intersectional bias and evaluation of group bias even when the number of samples is low.
Figure 1: An illustration of why evaluating marginal bias does not suffice. A classifier accepts the same proportion of each age and race group, but rejects the entire subgroup (group intersection) of young blue people. This bias would go undetected unless group intersections are considered. The graphic is a modification of Figure 1 in Němeček et al. (2025).
2 Challenges
We now outline several difficulties that arise when bias is estimated.
2.1 Sample Complexity and Computability
To estimate data quality qualtiy, i.e., the preexisting bias in the data per se, one generally calculates a distance between the distribution of values in the training data and the general population. There are many such distances on measure spaces: total variation, MMD (Gretton et al., 2012), and Wasserstein (Vaserstein, 1969), to name the most popular. However, the sample complexity of these distances varies; that is, the number of samples required to achieve an accurate approximation of the selected distance is very large. Moreover, some of them (esp. Wasserstein-1) are believed to be inapproximable (cf. Lee et al. (2023, 2024)).
2.2 Choice of Protected Attributes and Intersectional Aspects
The sample complexity is also clearly tied to the number of protected subgroups. In the case of studying the data quality, one may need to test the distance between the distribution of values in the training data and the general population for each protected subgroup. The number of protected subgroups may thus enter linearly in the overall sample complexity, while being exponential in the number of protected attributes.
For example, consider the The California Consumer Privacy Act, which lists a number of possible protected attributes and their values to be protected. Thus, it defines 5,772,124,799 protected subgroups (for the population of California, which is approximately fourty million) by some estimates (Němeček et al., 2025). Estimating the bias for each such protected subgroup is clearly challenging.
2.3 High-Dimensional Probability
Another potential issue comes from the fact that in a high-dimensional scenario, i.e., when the ambient dimension counts both protected and unprotected attributes, two points from a finite region, sampled uniformly at random, have distances that are increasingly concentrated around the mean Aggarwal et al. (2001). Also, considering the encoding of the dataset, be it one-hot or one-cold, there will always be a ”zero” value/field, which accounts for the people who refuse to provide a reply in the datasets coming from surveys. This added complexity is required even if the vast majority will respond accordingly.
2.4 Uncertainty Quantification
Obviously, any of the estimates of bias is meaningless without knowing the uncertainty in the estimate. If, for example, one learns that some measure of bias is estimated at 42, one also needs to know whether the error of the estimate is bounded from above by 1, 42 or even a higher value. If the error in the estimate is 42, there may be no bias at all. Estimating the uncertainty for non-linear functions, such as the bias measures above, is non-trivial. There are now first attempts (Němeček et al., 2025) to provide uncertainty quantification for such bias measures.
3 Solution
We propose a toolkit that implements two approaches to bias estimation. These have an advantage over previously considered approaches, namely the fact that the number of samples required is linear in the number of protected attributes, and the sample complexity can be controlled by controlling the error (or uncertainty) in the estimate.
Maximum Subgroup Discrepancy
When detecting bias, one has to look at all intersections of protected groups, because marginal fairness might be achieved despite exhibiting severe bias when looking at the intersections. As a type of curse of dimensionality, the number of group intersections (subgroups) grows exponentially with the number of protected attributes. This complicates the evaluation of intersectional bias in multiple ways. One has to face the exponential time complexity of testing each subgroup, while having exponentially fewer and fewer samples. A useful approach to this would be to consider the bias as a distance between two distributions, but as mentioned, most distributional distances have exponential sample complexity.
Maximum Subgroup Discrepancy (MSD) (Němeček et al., 2025) addresses the issues of high sample complexity of intersectional fairness. It is a distributional distance, defined as
$\mathrm{MSD}(\mu,\nu)=\sup_{S\in\mathcal{S}}|\mu(S)-\nu(S)|,$
where $\mu$ and $\nu$ are two distributions, $\mathcal{S}$ is the set of all subgroups, and $\mu(S)$ is the probability of sampling a point from the subgroup $S$ given the distribution $\mu$. In simpler terms, MSD looks for the subgroup with the highest difference in probability between two distributions. The MSD algorithm has linear sample complexity in the number of protected attributes, returns naturally defined subgroups (as conjunctions of feature value pairs), and the group that is guaranteed to have the highest discrepancy for practical dataset sizes with many protected attributes. Using mixed-integer optimization, it obtains the provably optimal solution (i.e. the maximally discrepant subgroup).
Subsampled Distances on Measure Spaces
Our second approach addresses the same issue regarding protected group intersections in a different manner. In particular, we subsample a test whether the distance between the two distributions is less than a given value ($\Delta$). For a given subgroup, we can check whether the given subgroup faces a bias within the threshold $\Delta$. In large datasets, we may have to subsample this test, but M. Matilla and Mareček (2025) have developed guaranties on the size of the error with a certain probability. Such guaranties are sometimes known as probably approximately correct (PAC) learning.
Usage of the toolkit
Figure 2: humancompatible.detect Workflow for bias evaluation
Using the humancompatible.detect toolkit is straightforward. The toolkit supports two common tasks in three input modes. The tasks are:
  • Detection of the most biased subgroup (according to MSD): returns a simple conjunctive rule (e.g., ”AGE between 18-25 AND SEX = Male”).
  • Quantification of a given subgroup: evaluate a given subgroup, or check. By selecting a different method:
    • MSD: Measuring the subgroup discrepancy of a given subgroup, with the input as the keyword argument rule.
    • l_inf: The pass/fail test against a tolerance, for which it requires both the attribute (feature_involved) and the particular value (subgroup_to_check) that will be analysed.
Regarding the input, one can utilize:
  • DataFrames: Detection occurs with the most_biased_subgroup(X, y); and evaluating with evaluate_biased_subgroup(X, y, …).
  • CSV: Detect with most_biased_subgroup_csv(csv_path, target_col); evaluate with evaluate_biased_subgroup_csv(csv_path, target_col, …).
  • Two-sample comparison: Detect with most_biased_subgroup_two_samples(X1, X2); evaluate with evaluate_biased_subgroup_two_samples(X1, X2, …).
    This mode finds which subgroup’s representation differs most between the two datasets. Both datasets must share the same columns.
4 Demonstration
We illustrate the toolkit on two examples, and show how users can (i) discover the most impacted subgroup using MSD, and (ii) check a concrete subgroup against a tolerance using the subsampled $\ell_{\infty}$ test.
4.1 Illustrative example
There are three columns: Race (categorical: Green, Blue, Purple), Age (categorical buckets: 0-18, 18-30, 30-45, 45-60), and a binary Target (e.g., approval). The dataset is built so that overall demographic parity holds, but one intersectional subgroup is disadvantaged, as shown in Figure 1.
  • We treat Race and Age as protected attributes, and use MSD to find the subgroup whose outcome rate differs most from the remainder of the population.
    The toolkit returns a rule as follows (after interpreting the rule):
    Race = Blue AND Age = 0-18
    After evaluating the bias for the obtained subgroup, we got the MSD value of $0.111\ldots$ (e.g., an 11.1% gap in positive-outcome rates between the subgroup and its complement).
    MSD confirms that young Blue individuals are the most disadvantaged group, even though the data looks fair on the margins. This is why checking intersections matters.
  • The new protected attribute Gender (categorical: M, F) was added to the previous dataset, while keeping the original demographic parity properties.
    We consider the concrete subgroup Gender = F, and test whether its distribution of positive outcomes differs from the full population by at most a tolerance $\delta$ in the $\ell_{\infty}$ norm.
    For a police threshold $\delta=0.125$, the method reports that the most impacted subgroup bias is greater than 0.125 (fail). In other words, the maximum difference between the subgroup and the overall positive outcome histograms is over the threshold.
Whereas MSD discovers the worst offending rule, the $\ell_{\infty}$ test validates a particular subgroup against a clear pass/fail standard.
4.2 Real-world ACS data
We also demonstrate the toolkit on the ACS Income task (folktables) using one state at a time, and a pair of states for comparison. We focus on a small set of protected attributes commonly used in practice (e.g., Age, Educational attainment, Occupation, Sex, Race), and a binary outcome (income >> \$50,000).
  • In the within-state case, where we want to discover the worst subgroup, we ran MSD separately for two states (Florida and New Hampshire). In both cases, we obtained the subgroup, and after evaluating its discrepancy:
    • Florida (FL). Rule: Race = White AND Sex = Male. MSD = 0.193. Interpreting the sign with a subsampled signed gap, $\Delta\approx+0.193$, this subgroup is over-represented among the positive class (high income) in FL.
    • New Hampshire (NH). Rule: Race = White AND Sex = Male. MSD = 0.217. With $\Delta\approx+0.217$, the subgroup is again over-represented among high-income individuals in NH.
  • For the cross-state comparison, we search for the subgroup where two states differ most. Compare two states directly (FL vs. NH) using all the protected attributes that are available for this survey. MSD finds the subgroup whose prevalence differs most across the two datasets.
    Rule: Race = White AND Citizenship = Born in the US AND Cognitive difficulty = No AND Difficulty hearing = No. With MSD = 0.226. The signed gap $\Delta\approx+0.226$ indicates this subgroup is more prevalent in NH than in FL.
This humancompatible.detect toolkit lets you go from tabular data to a concrete, human-readable finding in a few lines. For readers who wants to go through all the experiments, the accompanying notebooks are fully reproducible and annotated with small tips and additional explanations, so you can adapt them to your setting without going into the source code.
5 Conclusions
We have presented a toolkit for bias detection. Crucially, it can consider protected group intersections and provide bias evaluations with guarantees even for smaller datasets. It has an intuitive API and is designed for easy use by practitioners without requiring domain expertise.
Acknowledgments and Disclosure of Funding
This work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101070568 (AutoFair).