### v2

# humancompatible.detect: a Python Toolkit for Detecting Bias in AI Models

Abstract
There is a strong recent emphasis on trustworthy AI. In particular, international regulations, such as the AI Act, demand that AI practitioners measure data quality on the input and estimate bias on the output of high-risk AI systems. However, there are many challenges involved, including scalability (MMD) and computability (Wasserstein-1) issues of traditional methods for estimating distances on measure spaces.
Here, we present humancompatible.detect, a toolkit for bias detection that addresses these challenges. It incorporates two newly developed methods to detect and evaluate bias: maximum subgroup discrepancy (MSD) and subsampled $\ell_{\infty}$ distances. It has an easy-to-use API documented with multiple examples. humancompatible.detect is licensed under the Apache License, Version 2.0.
keywords:
artificial intelligence , fairness , bias detection , toolkit
2020 MSC:
68T01
journal: SoftwareX
Nr. Code metadata description Metadata
C1 Current code version 0.1.4
C2 Permanent link to code/repository used for this code version https://github.com/humancompatible/detect/tree/6b3fb21a155d012cb244c9b5b7ead1f8a978038f
C3 Permanent link to Reproducible Capsule none
C4 Legal Code License Apache-2.0
C5 Code versioning system used git
C6 Software code languages, tools, and services used Python, pandas, Pyomo, HiGHS, SciPy
C7 Compilation requirements, operating environments & dependencies python $\geq$ 3.10, pandas $\geq$ 2.2 pyomo $\geq$ 6.9, highspy $\geq$ 1.10, scipy $\geq$ 1.16
C8 If available Link to developer documentation/manual https://humancompatible-detect.readthedocs.io/en/latest/
C9 Support email for questions martige1@fel.cvut.cz
Table 1: Code metadata (mandatory)
1 Motivation and significance
With the increase in practical applications using machine-learning models and AI systems, new ethical problems have appeared. One particular problem is the proliferation (or even strengthening) of societal bias. Infamous examples of AI bias include hiring algorithms [5] or recidivism predictors [9].
In response, more and more regulatory frameworks [e.g., 20] and related standards [e.g., 8, 22] require testing for bias, without any clear agreement on the precise definition of said bias. Indeed, there are many bias measures [17].
The lack of a single definition of bias is understandable, considering the wide range of applications of AI systems. Non-exhaustively:
  • 1.
    Fairness in (large) language and/or vision models, the currently dominant AI models [e.g., 1, 4].
  • 2.
    Fairness in rankings: the main related case being job recruitment. See, for instance, [19, 11, 6].
  • 3.
    Fairness in recommender systems: such as the ones widely used currently by, e.g., YouTube or Spotify [19, 15].
  • 4.
    Spatial aspects of fairness: In contrast to continuous attributes, such as age and income, the case of discrimination based on location demands special attention [21, 12].
  • 5.
    Fairness for forecasting: Consider the well-known COMPAS dataset. It is traditionally treated as a classification task. A forecasting-based approach, however, obtains the best results currently reported on the dataset [29].
  • 6.
    Fairness in sharing-economy business models: such as the model of companies such as Uber and Cabify [28].
Even when agreeing on a particular definition of bias, several difficulties remain:
Sample Complexity and Computability
To estimate data quality, i.e., the preexisting bias in the data per se, one generally calculates a distance between the distribution of values in the training data and the general population. There are many such distances on measure spaces: total variation, MMD [7], and Wasserstein [26], to name the most popular. However, the sample complexity of these distances varies; that is, the number of samples required to achieve an accurate approximation of the selected distance is very large. Moreover, some of them (esp. Wasserstein-1) are believed to be inapproximable (cf. [13, 14]).
Framework Detection Intersections Low sample complexity Optimal Ref
Themis-ml - - [2]
FAT Forensics - - [24]
Fairlearn - - [27]
AI Fairness 360 ? [3]
FairTest [25]
humancompatible.detect here
Table 2: Comparison to common Python fairness frameworks in four dimensions. Whether they allow for detecting bias (not just measuring bias of a predefined group), if they allow for considering group intersections natively, whether some of those detection methods have low sample complexity, and whether the methods have (optimality) guarantees. When no detection method is available, we omit the last two dimensions. The most similar, FairTest, focuses on identifying protected subgroups using a tree structure that aligns with the predictor’s output. Importantly, none of the other frameworks consider the MSD distance or the subsampled $\ell_{\infty}$ norm.
Choice of Protected Attributes and Intersectional Aspects
The sample complexity is also clearly tied to the number of protected subgroups. In the case of studying the data quality, one may need to test the distance between the distribution of values in the training data and the general population for each protected subgroup. The number of protected subgroups may thus enter linearly in the overall sample complexity, while being exponential in the number of protected attributes.
For example, consider The California Consumer Privacy Act, which lists a number of possible protected attributes and their values to be protected. The number of protected subgroups can go up to 5,772,124,799 on US Census data [18]. Estimating the bias for each such protected subgroup is clearly challenging.
Other fairness toolkits
There is a variety of existing toolkits, of which only some allow for bias detection. Overall are mostly concerned with either training fair models or fairness evaluation. We provide an overview of various toolkits in Table 2. There are additional toolkits focused on visualiation (e.g., What-If Tool), however, we ommit those from the comparison since they are visual tools rather than evaluation toolkits.
Themis-ml [2] and Fairlearn [27] do not allow for detecting (most) biased subgroups, and detection part of FAT Forensics [24] does not allow for the study of intersections of protected groups. AI Fairness 360 [3] contains only the MDSS method for the detection of intersectional bias, which uses a non-standard definition of bias and it’s sample complexity is unclear. It is also a probabilistic method. FairTest [25] contains methods with low sample complexity, however these only test how aligned are protected attributes to the prediction, evaluating correlations rather than standard bias measures. They form a greedy tree, separating groups into intersectional subgroups, without guarantees on optimality.
2 Software description
Figure 1: A diagram of possible uses of the humancompatible.detect toolkit. In addition to evaluating bias of a protected group, it allows for detection of the most biased subgroup given the dataset.
We introduce humancompatible.detect, an open-source toolkit of methods for bias detection. Designed for ease of use, its simple API and guided examples are accessible to non-expert users. It is available on PyPI [^1] https://pypi.org/project/humancompatible-detect/. The toolkit allows for reliable detection of intersectional bias and evaluation of group bias even when the number of samples is low. Fig 1 illustrates the use of the toolkit.
2.1 Software architecture
As per Figure 1, the two main toolkit objectives are detect and evaluate. The first one spots the most biased subgroup in the data, while the second informs us about the bias against a given subgroup within the dataset. The data can be passed in multiple ways.
Input data
There are three modes of input:
  • 1.
    In-memory tables (pandas DataFrame). The dataset is passed in as the input data X and the binary target column y.
  • 2.
    CSV file, with a designated target column. Requiring the path to this data and the name of the column.
  • 3.
    Two-sample comparison, where the goal is to identify subgroups whose representations differs most between two distributions, represented by the datasets. Both datasets must share the same columns. One simply passes two pandas DataFrames X1 and X2 as arguments.
Detect bias
Figure 2: detect_bias activity diagram. Depending on the input mode used, different functions can be run, always calling most_biased_subgroup. After that, the data preprocessing takes places with prepare_dataset. Subsequently, the program stops if a method different from MSD has been selected, otherwise outputs the most affected subgroup as a result. This allows for easy extensibility. More details on the prepare_dataset are in Figure 4.
To detect bias in data the toolkit alows to use MSD. It can check the exponentially growing number of dataset subgroups and spot the most biased subgroup in the dataset. It returns a simple conjunctive rule representation of a subgroup (e.g.,“AGE between 18-25 AND SEX = Male”). The toolkit can be easily extended to allow for other methods, such as SPSF [10]. See Figure 2 for a graphical description.
Evaluate bias
Figure 3: evaluate_bias activity diagram. Both input and preprocessing are identical to detect_bias, with the addition of keyword arguments specifying the (sub)group (Group definition) and potentially other arguments (e.g., the $\Delta$ threshold). The output can either be a value (e.g., subgroup discrepancy) or a boolean value whether bias is above a given threshold (e.g., in subsampled $\ell_{\infty}$). More details on the prepare_dataset are in Figure 4.
Independent of learning which subgroup is the most (dis)advantaged within the data, we can quantify the bias for any subgroup. The MSD method measures the subgroup discrepancy of a given subgroup, which is specified as a conjunction of attributes. Additionally, with our second method, subsampled $\ell_{\infty}$, we can test whether the bias in a particular subgroup is greater or smaller than a value of our choice. For this task, both methods require additional parameters to be passed as keyword arguments. See Figure 3 for a graphical visualization of the process.
Internally, both tasks use a routine to preprocess the input data, prepare_dataset, which is also illustrated in further detail in Fig 4.
Figure 4: prepare_dataset activity diagram. This function preprocesses the input data in the form of a pandas DataFrame, y, being the target column and X, the rest of the data. It prunes incomplete rows, allows for binarization, and also for specialized feature processing, such as mapping categorical variables into variables with fewer distinct values or custom binning of continuous variables. It also allows for subsampling the data, if passed data is too large. This can speed up the processing without reducing accuracy too much, since implemented methods have low sample complexity.
For convenience, we provide a unified wrapper for users, detect_and_score(…) combining the detection and evaluation into one step. It takes the data (in either of the three formats), the protected attributes and target specification, and returns the corresponding subgroup description together with the computed bias score (or pass/fail decision, depending on the method).
2.2 Software functionalities
We propose a toolkit that implements two approaches to bias estimation. With these, users can (i) discover the most impacted subgroup using Maximum Subgroup Discrepancy, and (ii) check a concrete subgroup against a tolerance using the subsampled $\ell_{\infty}$ test.
Maximum Subgroup Discrepancy
When detecting bias, it is crucial to look at all intersections of protected groups, because marginal fairness might be achieved despite exhibiting severe bias when looking at the intersections. As a type of curse of dimensionality, the number of group intersections (subgroups) grows exponentially with the number of protected attributes. This complicates the evaluation of intersectional bias in multiple ways. One has to face the exponential time complexity of testing each subgroup, while having exponentially fewer and fewer samples. A useful approach to this would be to consider the bias as a distance between two distributions, but as mentioned, most distributional distances have exponential sample complexity.
Maximum Subgroup Discrepancy (MSD) [18] addresses the issues of high sample complexity of intersectional fairness. It is a distributional distance, defined as
$\mathrm{MSD}(\mu,\nu)=\sup_{S\in\mathcal{S}}|\mu(S)-\nu(S)|,$
where $\mu$ and $\nu$ are two distributions, $\mathcal{S}$ is the set of all subgroups, and $\mu(S)$ is the probability of sampling a point from the subgroup $S$ given the distribution $\mu$. In simpler terms, MSD looks for the subgroup with the highest difference in probability between two distributions. The MSD algorithm has linear sample complexity in the number of protected attributes, returns naturally defined subgroups (as conjunctions of feature value pairs), and the group that is guaranteed to have the highest discrepancy for practical dataset sizes with many protected attributes. Using mixed-integer optimization, it obtains the provably optimal solution (i.e. the maximally discrepant subgroup).
Subsampled Distances on Measure Spaces
Our second approach addresses the same issue regarding protected group intersections in a different manner. In particular, we subsample a test whether the distance between the two distributions is less than a given value ($\Delta$). For a given subgroup, we can check whether the given subgroup faces a bias within the threshold $\Delta$. In large datasets, we may have to subsample this test, but [16] have developed guaranties on the size of the error with a certain probability. Such guaranties are sometimes known as probably approximately correct (PAC) learning.
3 Illustrative example
We illustrate the toolkit on a synthetic example dataset.
There are three columns: Race (categorical: Green, Blue, Purple), Age (categorical buckets: 0-18, 18-30, 30-45, 45-60), and a binary Target (e.g., approval). The dataset is built so that marginal fairness holds, but one intersectional subgroup is disadvantaged, as shown in Figure 5.
We treat Race and Age as protected attributes, and use MSD to find the subgroup.
from humancompatible.detect import detect_and_score
rule, msd_val = detect_and_score(
csv_path = "01_data.csv",
target_col = "Target",
protected_list = ["Race", "Age"],
method = "MSD",
The toolkit returns a rule as follows (after interpreting the rule):
print("The group is: " +
" AND ".join(str(r) for _, r in rule))
print("MSD value is:", msd_val)
The group is: Race = Blue AND Age = 0-18
MSD value is: 0.111111
After evaluating the bias for the obtained subgroup, we got the MSD value of $0.111\ldots$ (e.g., an 11.1% gap between the positive-outcome and the negative outcome for the given subgroup). MSD confirms that young Blue individuals are the most disadvantaged group, even though the data looks fair on the marginals. This is why checking intersections matters.
To test the subsampled $\ell_{\infty}$ norm, we add a new protected attribute Gender (categorical: M, F) the previous dataset, while keeping the original demographic parity properties. We consider the concrete subgroup Gender = F, and test whether its distribution of positive outcomes differs from the full population by at most a tolerance $\Delta$ in the $\ell_{\infty}$ norm.
For a threshold $\Delta=0.125$, the method reports that the most impacted subgroup bias is greater than 0.125 (fail). In other words, the maximum difference between the subgroup and the overall positive outcome histograms is over the threshold.
4 Impact
Figure 5: An illustration of why evaluating marginal bias does not suffice. A classifier accepts the same proportion of each age and race group, but rejects the entire subgroup (group intersection) of young blue people. This bias would go undetected unless group intersections are considered. The graphic is a modification of Figure 1 in [18].
Even in the best-studied case of fairness of classification, most of the proposed bias measures, such as statistical parity or equality of opportunity, focus on protected groups in isolation. This approach often misses the intersections of groups or requires one to sift through an exponential number of group intersections (subgroups). Considering all subgroups is crucial, as illustrated by Figure 5, where data is marginally fair, but intersectionally biased.
Another view of bias evaluation in cases when many subgroups need to be considered is bias detection. That is, detecting where bias exists in the data, rather than measuring it entirely over all protected features. Detecting such a biased subgroup allows focusing on the particular subgroup [e.g., 23]. This way, one can address the most significant biases individually through mitigation techniques tailored to the particular group, instead of treating all biases the same.
We showed that humancompatible.detect addresses both questions (intersectional fairness and bias detection) by allowing its users to go from tabular data to a concrete, human-readable finding in a few lines. Moreover, its two implemented methods, have an advantage over previously considered approaches such as those based on MMD [7] and Wasserstein [26] distances, in terms of the number of samples required. For MSD, the number of samples is linear in the number of protected attributes, and the sample complexity can be controlled by controlling the error (or uncertainty) in the estimate.
5 Conclusions
We have presented a toolkit for bias detection. Crucially, it can consider protected group intersections and provide bias evaluations with guarantees even for smaller datasets. It has an intuitive API and is designed for easy use by practitioners without requiring domain expertise.
CRediT author statement
Germán M. Matilla: Software, Visualization, Validation, Writing - Original Draft, Jiří Němeček: Methodology, Software, Writing - Original Draft, Visualization, Writing - Review & Editing, Illia Kryvoviaz: Software, Validation, Visualization, Writing - Original Draft, Jakub Mareček: Conceptualization, Supervision, Writing - Original Draft, Writing - Review & Editing.